# biblio.bib

### v1

@@ 4 Algorithms for the Critical Values
It gives as output an approximation of the solution to ($\mathcal{H}$JE) with initial datum $\phi$, which we denote by $v^{\Delta}$.
The domain of $v^{\Delta}$ is a grid $\Gamma^{T}_{\Delta}$ of $\Gamma\times[0,T]$ defined in the following way: we fix an orientation $\mathcal{E}^{+}$, i.e., a subset of $\mathcal{E}$ containing exactly one arc in each pair $\{\gamma,\widetilde{\gamma}\}$, and set
@@
[…]
@@
We then introduce the grids
$\mathscr{S}_{\Delta,\gamma}\coloneqq\left\{s_{i}^{\Delta,\gamma}:i\in\left\{0,\dotsc,N_{\gamma}^{\Delta}\right\}\right\},\quad\text{for }\gamma\in\mathcal{E}^{+},\qquad\mathscr{T}_{\Delta}^{T}\coloneqq\left\{t^{\Delta}_{m}:m\in\left\{0,\dotsc,N^{\Delta}_{T}\right\}\right\}$
and finally
@@
$0<\Delta x<|\gamma|,\quad\text{for every }\gamma\in\mathcal{E},\qquad 0<\Delta t<T\qquad\text{and}\qquad\Delta t\leq\min_{\gamma\in\mathcal{E}^{+}}\frac{\Delta_{\gamma}x}{\beta_{0}},$ (14)
where $\beta_{0}$ is a constant depending on the Hamiltonian $\mathcal{H}$ and the initial datum $\phi$ defined in [CarliniSiconolfi23].
@@
[CarliniCoscettiPozza24, Corollary 4.14] Fixed a Lipschitz continuous initial datum $\phi$, $T>0$ and an admissible pair $\Delta$, let $v^{\Delta}$ be the approximation of $\mathcal{S}(t)\phi$ on $\Gamma^{T}_{\Delta}$ given by the semi-Lagrangian scheme described in [CarliniSiconolfi23]. There exists a constant $C>0$, independent of $\Delta$ and $T$, such that
@@
We stress out that, in principle, our algorithms could employ any other numerical method for the approximation of solutions to ($\mathcal{H}$JE). In view of [CarliniCoscettiPozza24, Theorem 3.4], which proves that our definition of solution is equivalent to the one proposed in [ImbertMonneau17], we could even choose the finite difference schemes proposed in [CamilliCarliniMarchi18, CostesequeLebacqueMonneau14] or the semi-Lagrangian method outlined in [CarliniFestaForcadel20].
@@ 4.1 An Algorithm with an A Priori Error Estimate
Using an approach employed in [AchdouCamilliCapuzzo-Dolcetta08] for equations posed on the $N$-dimensional torus, we obtain an algorithm complete with an error estimate which proves the convergence of the scheme. The main idea behind this algorithm is to exploit the estimates given in Propositions 3.12 and 3.13.
Algorithm 1 consists of two moduli: $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$. $\mathcal{M}_{1}$ takes as inputs a Lipschitz continuous initial datum $\phi$, a final time $T>0$ and an admissible pair $\Delta\coloneqq(\Delta x,\Delta t)$, and gives as output an approximation $v^{\Delta}$ of the exact solution to ($\mathcal{H}$JE) with initial datum $\phi$, presented at the start of this Section, at time $T$:
@@
In view of Lemma 3.13 and (4), Algorithm 1 can be improved by taking
@@
$\left|\frac{\overline{c}^{\Delta}_{k}+\underline{c}^{\Delta}_{k}}{2}-c\right|\leq C_{0}\left(\frac{1+\Delta x}{kT}+\frac{1}{\sqrt{\Delta t}}\min\left\{\Delta x,\frac{\Delta x^{2}}{\Delta t}\right\}+\sqrt{\Delta t}\right),\qquad\text{for any }k\geq 1,$ (15)
@@
4.4 proves that the approximation given by Algorithm 1 converges to the critical value as $k$ diverges and, if (14) holds,
@@
Remark 4.5.
While Theorem 4.4 provides an a priori error estimate, it is not feasible. Indeed, even ignoring the error of the scheme $v^{\Delta}$ and assuming $\operatorname{diam}(\Gamma)\approx 1$, the presence of $1/kT$ in (15) implies that to obtain an approximation of the critical value with a tolerance of $10^{-3}$ this estimate would require a total time of approximation $kT\approx 10^{3}$, which is much bigger than the ones seen in the simulations in Section 5. We point out that, for a reasonable pair $\Delta$, a high value of $kT$ correspond to a high number of iterations, if $T$ is small, or a high computational cost for $v^{\Delta}$.
Proof of Theorem 4.4.
@@
where $\ell$ is the Lipschitz constant of $(x,t)\mapsto(\mathcal{S}(t)\phi)(x)$. Then, by Theorem 4.1, there is $C>0$ independent of $\Delta$ and $k$ such that
@@
thus Proposition 3.12 yields
@@
Finally, we know from Lemma 3.13 that, for any $k\geq 1$,
@@
thus, arguing as in the proof of Theorem 4.4, we get the following Proposition which justifies the stopping criterion $\overline{c}^{\Delta}_{k}-\underline{c}^{\Delta}_{k}<2\varepsilon$ at 8 in Algorithm 1.
@@
Under the same assumptions of Theorem 4.4, we have that
@@ 4.2 A Faster Iterative Algorithm
Next we describe a numerical scheme which exploit Proposition 3.14, instead of Proposition 3.12, to obtain an approximation of the critical value. While we do not have an a priori error estimate for this new algorithm, the simulations in Section 5 show that it is considerably faster than Algorithm 1.
Following Lemma 3.15, $c_{k}$ defined in (13) is a reasonable approximation of $c$ for $k$ big enough, thus we propose a scheme, Algorithm 2, which estimates the critical value using a numerical approximation of $c_{k}$. It consists of the same two moduli $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ used in Algorithm 1. They are used to defining at each iteration, fixed a Lipschitz continuous initial datum $\phi$, a time $T>0$ and an admissible pair $\Delta$,
@@
In view of Proposition 3.14 and (4), Algorithm 2 can be improved by taking
@@
The next Theorem shows the convergence of Algorithm 2.
@@
and $c_{k}$ as in Lemma 3.15, we have that,
@@
where $\ell$ is the Lipschitz constant of $(x,t)\mapsto(\mathcal{S}(t)\phi)(x)$. Setting $\widetilde{c}_{k}$ and $\undertilde{c}_{k}$ as in Proposition 3.14, Theorem 4.1 yields that there is a $C>0$ independent of $\Delta$ and $k$ such that
@@
This and Lemma 3.15 imply (17) whenever (14) and (18) hold. ∎
Adopting the same notation used in the proof of Theorem 4.8, we notice that Lemma 3.15 yields
@@
This and Theorem 4.8 prove the following error estimate for Algorithm 2.
@@
Under the same assumptions of Theorem 4.8, we have that, for any $k\geq 1$,
@@
4.9 justifies the stopping criterion $\widetilde{c}^{\Delta}_{k}-\undertilde{c}^{\Delta}_{k}<2\varepsilon$ at 8 in Algorithm 2, but does not prove the convergence of the algorithm, as Theorem 4.8, due to the presence of the terms $\widetilde{c}^{\Delta}_{k}$ and $\undertilde{c}_{k}^{\Delta}$.
@@
We point out that the presence of $k$ in the numerator of both (18) and (19) affects the estimated convergence rate of our method. It also means that the approximation provided by Algorithm 2 becomes less reliable with each step, in contrast with Algorithm 1. Fixed an admissible pair $\Delta$, one could exploit (19) to pick a maximal number of iterations after which the algorithm should not be trusted, namely, fix a maximal number of iterations Algorithm 2 can perform reliably.
@@ 5 Numerical Simulations
We now focus on the selection of parameters required by Algorithm 1. Based on the theoretical results established in the previous sections, we examine the expected performance of the algorithm as a function of these chosen parameters. First, we select the pair $\Delta$ such that
@@
If $\Delta t$ is further chosen to satisfy (14), then Theorem 4.4 ensures that for $k$ large enough the error $\left|\frac{\overline{c}^{\Delta}_{k}+\underline{c}^{\Delta}_{k}}{2}-c\right|$ is bounded above by a term proportional to $\min\left\{\frac{\Delta x}{\sqrt{\Delta t}},\frac{\Delta x^{2}}{\Delta t^{3/2}}\right\}$ and independent of $k$. This term arises solely from the error estimate for the semi-Lagrangian scheme employed to approximate the solution to ($\mathcal{H}$JE) (as shown in Theorem 4.1), it remains constant for fixed $\Delta$ and vanishes as $\Delta\to 0$. Equivalently, by selecting a small enough stopping condition $\varepsilon$ in Algorithm 1, Proposition 4.6 implies that the error $\left|\frac{\overline{c}^{\Delta}_{k}+\underline{c}^{\Delta}_{k}}{2}-c\right|$ is bounded above by a term proportional to $\min\left\{\frac{\Delta x}{\sqrt{\Delta t}}\frac{\Delta x^{2}}{\Delta t^{3/2}}\right\}$ when the algorithm reaches this tolerance. As before, this term remains constant for fixed $\Delta$ and tends to zero as $\Delta\to 0$. Consequently, this analysis enables us to conclude that, for a fixed $\Delta$, the tolerance $\varepsilon$ can be chosen as a function of $\Delta$ without any loss of accuracy.
@@
In the next simulations we implement the algorithm with time step $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta x_{\gamma})/\beta_{0}$, for which both (20) and (14) hold. We also evaluate the algorithm for $\Delta t=\Delta x/2$ and $\Delta t=\Delta x^{5/6}$, which satisfy (20) but do not satisfy (14). These last two choices do not meet the assumptions in Theorem 4.1 and, consequentially, in Theorem 4.4, but allow us for larger time steps. However, the last choice aligns with the convergence properties of the semi-Lagrangian scheme for the time dependent Hamilton–Jacobi equation ($\mathcal{H}$JE) established in [CarliniSiconolfi23].
@@
We additionally perform these tests for Algorithm 2 to compare its performance with Algorithm 1. We point out that, while Theorem 4.4 guarantees convergence of Algorithm 1 as $\Delta\to 0$ suitably and as $k\to\infty$ independently of $\Delta$, the convergence of Algorithm 2 is proven if (18) holds. Moreover, while Proposition 4.6 ensures that, by choosing a sufficiently small stopping condition, the difference between the approximated and exact critical value is bounded above by a term dependent on $\Delta$ and independent of $k$ when Algorithm 1 reaches the prescribed tolerance, Proposition 4.9 does not provide a similar result for Algorithm 2.
@@
More specifically, we implement Algorithms 1 and 2 with the improvements detailed in Remarks 4.3 and 4.7. With a slight abuse of notation, we denote by $c^{\Delta}$ the output returned by both Algorithm 1 and Algorithm 2. The accuracy is measured by computing the absolute value of the difference between the approximated critical value $c^{\Delta}$ and a reference critical value $\hat{c}$ or the exact critical value $c$, when it is known. We apply the algorithms in two ways: either by fixing the number of iterations to be executed or by setting a tolerance upon reaching which the algorithm terminates. In the following tables, the column labeled $k$ reports the number of iterations performed, either to complete the predetermined number of iterations or to satisfy the prescribed tolerance. In the following simulations, we run the algorithms with $\phi\equiv 0$, $T=1$ and minimal flux limiters.

### v2

@@ 4.1 Numerical Approximation of Time-Dependent HJ Equations
To define the space-time grid, we fix an orientation $\mathcal{E}^{+}$, i.e., a subset of $\mathcal{E}$ containing exactly one arc in each pair $\{\gamma,\widetilde{\gamma}\}$, and set
@@
where $\lceil\cdot\rceil$ stands for the ceiling function. We define
@@
We introduce the grids associated to any $\gamma\in\mathcal{E}^{+}$ and the grid on the interval $[0,T]$
$\mathscr{S}_{\Delta,\gamma}\coloneqq\left\{s_{i}^{\Delta,\gamma}:i\in\left\{0,\dotsc,N_{\gamma}^{\Delta}\right\}\right\},\quad\text{for }\gamma\in\mathcal{E}^{+},\qquad\mathscr{T}_{\Delta}^{T}\coloneqq\left\{t^{\Delta}_{m}:m\in\left\{0,\dotsc,N^{\Delta}_{T}\right\}\right\},$
and finally the grid on $\Gamma$ and the grid on $\Gamma\times[0,T]$
@@
$0<\Delta x<|\gamma|,\quad\text{for every }\gamma\in\mathcal{E},\qquad 0<\Delta t<T\qquad\text{and}\qquad\Delta t\leq\min_{\gamma\in\mathcal{E}^{+}}\frac{\Delta_{\gamma}x}{\beta_{0}}.$ (14)
We denote by $B\left(\Gamma^{0}_{\Delta}\right)$ and $B(\mathscr{S}_{{}_{\Delta},\gamma})$, for $\gamma\in\mathcal{E}^{+}$, the spaces of bounded functions from $\Gamma^{0}_{\Delta}$ and $\mathscr{S}_{{}_{\Delta},\gamma}$, respectively, to $\mathds{R}$. Given an arc $\gamma\in\mathcal{E}^{+}$ and $f\in B(\mathscr{S}_{{}_{\Delta},\gamma})$, we define the operator
$S_{\gamma}[f](s)\coloneqq\min_{\frac{s-|\gamma|}{\Delta t}\leq\lambda\leq\frac{s}{\Delta t}}\{I_{\gamma}[f](s-\Delta t\lambda)+\Delta t\widetilde{L}_{\gamma}(s,\lambda)\},\qquad\text{for }s\in\mathscr{S}_{{}_{\Delta},\gamma},$
where $I_{\gamma}$ is the interpolating polynomial of degree 1 defined by
$I_{\gamma}[f](s)\coloneqq f(s_{i})+\frac{s-s_{i}}{s_{i+1}-s_{i}}(f(s_{i+1})-f(s_{i})),$
with $s_{i},s_{i+1}\in\mathscr{S}_{{}_{\Delta},\gamma}$ and $s\in[s_{i},s_{i+1}]$. To define the scheme, we extend $S_{\gamma}$ to $B\left(\Gamma^{0}_{\Delta}\right)$ setting the map $S:B\left(\Gamma^{0}_{\Delta}\right)\to B\left(\Gamma^{0}_{\Delta}\right)$ such that
  • if $x\in\Gamma^{0}_{\Delta}\setminus\mathbf{V}$ let $\gamma\in\mathcal{E}^{+}$ and $s\in\mathscr{S}_{{}_{\Delta},\gamma}$ be such that $x=\gamma(s)$, then
    $S[f](x)\coloneqq S_{\gamma}[f\circ\gamma](s);$
  • if $x\in\mathbf{V}$ we set $S$ through a two steps procedure:
    $\widetilde{S}[f](x)\coloneqq\,$ $\min\left\{S_{\gamma}[f]\left(\gamma^{-1}(x)\right):\text{$\gamma\in\mathcal{E}^{+}$ with $\gamma(0)=x$ or $\gamma(|\gamma|)=x$}\right\},$
    $S[f](x)\coloneqq\,$ $\min\left\{\widetilde{S}[f](x),f(x)-c_{x}\Delta t\right\}.$
Given a Lipschitz initial datum $\phi$, the scheme computes the approximate solution $v^{\Delta}:\Gamma^{T}_{\Delta}\to\mathds{R}$ to the time-dependent problem in the following way:
$\left\{\begin{aligned} v^{\Delta}(x,0)=\,&\phi(x),&&\text{if }x\in\Gamma^{0}_{\Delta},\\ v^{\Delta}(x,t)=\,&S\left[v^{\Delta}(\cdot,t-\Delta t)\right](x),&&\text{if }t\neq 0,\,(x,t)\in\Gamma^{T}_{\Delta}.\end{aligned}\right$. (Discr)
@@
[11, Corollary 4.14] Fixed a Lipschitz continuous initial datum $\phi$, $T>0$ and an admissible pair $\Delta$, let $v^{\Delta}$ be the approximation of $\mathcal{S}(t)\phi$ on $\Gamma^{T}_{\Delta}$ given by the semi-Lagrangian scheme (Discr). There exists a constant $C>0$, independent of $\Delta$ and $T$, such that
@@
We stress out that, in principle, our algorithms could employ any other numerical method for the approximation of solutions to ($\mathcal{H}$JE). In view of [11, Theorem 3.4], which proves that our definition of solution is equivalent to the one proposed in [17], we could even choose the finite difference schemes proposed in [6, 14] or the semi-Lagrangian method outlined in [10].
@@ 4.2 An Algorithm with an A Priori Error Estimate
Using an approach employed in [1] for equations posed on the $N$-dimensional torus, we obtain an algorithm complete with an error estimate which proves the convergence of the scheme. The main idea behind this algorithm is to exploit the estimates given in Theorems˜3.8 and 3.9.
Algorithm˜1 consists of two moduli: $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$. $\mathcal{M}_{1}$ is the scheme (Discr), takes as inputs a Lipschitz continuous initial datum $\phi$, a final time $T>0$ and an admissible pair $\Delta\coloneqq(\Delta x,\Delta t)$, and gives as output an approximation $v^{\Delta}$ of the exact solution to ($\mathcal{H}$JE) with initial datum $\phi$, presented at the start of this Section, at time $T$:
@@
In view of Corollary˜3.9 and (5), Algorithm˜1 can be improved by taking
@@
$\left|\frac{\overline{c}^{\Delta}_{k}+\underline{c}^{\Delta}_{k}}{2}-c\right|\leq C_{0}\left(\frac{1+\Delta x}{kT}+\frac{1}{\sqrt{\Delta t}}\min\left\{\Delta x,\frac{\Delta x^{2}}{\Delta t}\right\}+\sqrt{\Delta t}\right),\qquad\text{for any }k\geq 1,$
@@
˜4.4 proves that the approximation given by Algorithm˜1 converges to the critical value as $k$ diverges and, if (14) holds,
@@
Proof of Theorem˜4.4.
@@
where $\ell$ is the Lipschitz constant of $(x,t)\mapsto(\mathcal{S}(t)\phi)(x)$. Then, by Theorem˜4.1, there is $C>0$ independent of $\Delta$ and $k$ such that
@@
thus Theorem˜3.8 yields
@@
Finally, we know from Corollary˜3.9 that, for any $k\geq 1$,
@@
thus, arguing as in the proof of Theorem˜4.4, we get the following Proposition which justifies the stopping criterion $\overline{c}^{\Delta}_{k}-\underline{c}^{\Delta}_{k}<2\varepsilon$ at ˜8 in Algorithm˜1.
@@
Under the same assumptions of Theorem˜4.4, we have that
@@ 4.3 A Faster Iterative Algorithm
Next we describe a numerical scheme which exploit Theorem˜3.10, instead of Theorem˜3.8, to obtain an approximation of the critical value. While we do not have an a priori error estimate for this new algorithm, the simulations in Section˜5 show that it is considerably faster than Algorithm˜1.
Following Corollary˜3.11, $c_{k}$ defined in (13) is a reasonable approximation of $c$ for $k$ big enough, thus we propose a scheme, Algorithm˜2, which estimates the critical value using a numerical approximation of $c_{k}$. It consists of the same two moduli $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ used in Algorithm˜1. They are used to defining at each iteration, fixed a Lipschitz continuous initial datum $\phi$, a time $T>0$ and an admissible pair $\Delta$,
@@
In view of Theorem˜3.10 and (5), Algorithm˜2 can be improved by taking
@@
The next Theorem shows the convergence of Algorithm˜2.
@@
and $c_{k}$ as in Corollary˜3.11, we have that,
@@
where $\ell$ is the Lipschitz constant of $(x,t)\mapsto(\mathcal{S}(t)\phi)(x)$. Setting $\widetilde{c}_{k}$ and $\undertilde{c}_{k}$ as in Theorem˜3.10, Theorem˜4.1 yields that there is a $C>0$ independent of $\Delta$ and $k$ such that
@@
This and Corollary˜3.11 imply (16) whenever (14) and (17) hold. ∎
Adopting the same notation used in the proof of Theorem˜4.7, we notice that Corollary˜3.11 yields
@@
This and Theorem˜4.7 prove the following error estimate for Algorithm˜2.
@@
Under the same assumptions of Theorem˜4.7, we have that, for any $k\geq 1$,
@@
˜4.8 justifies the stopping criterion $\widetilde{c}^{\Delta}_{k}-\undertilde{c}^{\Delta}_{k}<2\varepsilon$ at ˜8 in Algorithm˜2, but does not prove the convergence of the algorithm, as Theorem˜4.7, due to the presence of the terms $\widetilde{c}^{\Delta}_{k}$ and $\undertilde{c}_{k}^{\Delta}$.
@@
We point out that the presence of $k$ in the numerator of both (17) and (18) affects the estimated convergence rate of our method. It also means that the approximation provided by Algorithm˜2 becomes less reliable with each step, in contrast with Algorithm˜1. Fixed an admissible pair $\Delta$, one could exploit (18) to pick a maximal number of iterations after which the algorithm should not be trusted, namely, fix a maximal number of iterations Algorithm˜2 can perform reliably.
@@ 5 Numerical Simulations
We now focus on the selection of parameters required by Algorithm˜1. Based on the theoretical results established in the previous sections, we examine the expected performance of the algorithm as a function of these chosen parameters. First, we select a sequence of pairs $\Delta$ such that
@@
If $\Delta t$ is further chosen to satisfy (14), then Theorem˜4.4 ensures that for $k$ large enough the error $\left|\frac{\overline{c}^{\Delta}_{k}+\underline{c}^{\Delta}_{k}}{2}-c\right|$ is bounded above by a term proportional to $\min\left\{\frac{\Delta x}{\sqrt{\Delta t}},\frac{\Delta x^{2}}{\Delta t^{3/2}}\right\}$ and independent of $k$. This term arises solely from the error estimate for the semi-Lagrangian scheme employed to approximate the solution to ($\mathcal{H}$JE) (as shown in Theorem˜4.1), it remains constant for fixed $\Delta$ and vanishes as $\Delta\to 0$. Equivalently, by selecting a small enough stopping condition $\varepsilon$ in Algorithm˜1, Proposition˜4.5 implies that the error $\left|\frac{\overline{c}^{\Delta}_{k}+\underline{c}^{\Delta}_{k}}{2}-c\right|$ is bounded above by a term proportional to $\min\left\{\frac{\Delta x}{\sqrt{\Delta t}},\frac{\Delta x^{2}}{\Delta t^{3/2}}\right\}$ when the algorithm reaches this tolerance. As before, this term remains constant for fixed $\Delta$ and tends to zero as $\Delta\to 0$. Consequently, this analysis enables us to conclude that, for a fixed $\Delta$, the tolerance $\varepsilon$ can be chosen as a function of $\Delta$ without any loss of accuracy.
@@
In the next simulations we implement the algorithm with time step $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta x_{\gamma})/\beta_{0}$, for which both (19) and (14) hold. We also evaluate the algorithm for $\Delta t=\Delta x/2$ and $\Delta t=\Delta x^{5/6}$, which satisfy (19) but do not satisfy (14). These last two choices do not meet the assumptions in Theorem˜4.1 and, consequentially, in Theorem˜4.4, but allow us for larger time steps. However, the last choice aligns with the convergence properties of the semi-Lagrangian scheme for the time-dependent Hamilton–Jacobi equation ($\mathcal{H}$JE) established in [12].
@@
We additionally perform these tests for Algorithm˜2 to compare its performance with Algorithm˜1. We point out that, while Theorem˜4.4 guarantees convergence of Algorithm˜1 as $\Delta\to 0$ suitably and as $k\to\infty$ independently of $\Delta$, the convergence of Algorithm˜2 is proven if (17) holds. Moreover, while Proposition˜4.5 ensures that, by choosing a sufficiently small stopping condition, the difference between the approximated and exact critical value is bounded above by a term dependent on $\Delta$ and independent of $k$ when Algorithm˜1 reaches the prescribed tolerance, Proposition˜4.8 does not provide a similar result for Algorithm˜2. However, from a numerical point of view, Algorithm˜2 turns out to be highly advantageous. Indeed, the following simulations show that it attains the same accuracy as Algorithm˜1 while reducing the number of iterations by about 81% on average.
@@
We implement Algorithms˜1 and 2 with the improvements detailed in Remarks˜4.3 and 4.6. With a slight abuse of notation, we denote by $c^{\Delta}$ the output returned by both Algorithm˜1 and Algorithm˜2. The accuracy is measured by computing the absolute value of the difference between the approximated critical value $c^{\Delta}$ and a reference critical value $\hat{c}$ or the exact critical value $c$, when it is known. We apply the algorithms in two ways: either by fixing the number of iterations to be executed or by setting a tolerance upon reaching which the algorithm terminates. In the following tables, the column labeled $k$ reports the number of iterations performed, either to complete the predetermined number of iterations or to satisfy the prescribed tolerance. In the following simulations, we run the algorithms with $\phi\equiv 0$, $T=1$ and minimal flux limiters.